### v4

# The Minimal Attached Eddy in Wall Turbulence: Statistical Foundations, Inverse Identification and Influence Kernels

Abstract
Townsend’s attached eddy hypothesis models the logarithmic region of high Reynolds number wall turbulence as a random superposition of wall-attached, geometrically self-similar eddies whose sizes obey a scale-invariant population law. Building on the statistical framework of woodcock2015statistical, the present work (i) poses an inverse problem to infer the ideal single-eddy contribution (influence) functions for the mean velocity and Reynolds stresses from DNS moments, (ii) uses these inferred kernels to guide a minimal Biot–Savart-consistent hairpin-type eddy built from Rankine vortex rods together with an inviscid image system, and (iii) introduces a spectral Influence kernel that maps a self-similar eddy footprint to its one-dimensional energy spectrum. The Influence-kernel viewpoint yields a transparent explanation for the emergence (and limitations) of the linear part of the energy spectrum, provides a clear scale-by-scale decomposition and helps rationalize why simple eddy templates can reproduce a broad set of log-layer statistics once the mean-flow anchoring is fixed.
Characterizing and predicting the turbulent flow in the vicinity of a wall continues to not only be a problem of scientific interest, but also one of great practical relevance. Significant knowledge and insight has been gained from the development of experimental diagnostics in the 1960s to the emergence of direct numerical simulations in the 1980s, sophisticated measurements in the 1990s, an emphasis on coherent structures in the 2000s and operator-based dynamical models over the past decade. Against this backdrop of increasingly sophisticated experimentation, computations and analyses, the attached eddy hypothesis - originally proposed more than 50 years ago by  townsend1976structure - continues to serve as a simple, yet effective theory to qualitatively and quantitatively describe structural and statistical aspects of turbulent boundary layers, particularly focusing on the logarithmic region. The attached eddy model (AEM) was given a firm mathematical footing by  perry1982mechanism, and further developed by Marusic and co-workers over the past 25 years  (perry1995wall; marusic2013logarithmic; woodcock2015statistical; de2016influence). Among other useful features, the AEM is able to explain scaling behaviors of velocity moments, provide an explanation for uniform momentum zones (de2016uniform), and serve as a predictive model for the von-Karman constant as a function of Reynolds number. A review of the theory and developments of the AEM of wall turbulence can be found in  marusic2019attached. Over the past few years, theoretical and numerical analyses (e.g.  mckeon2019self; lozano2019characteristic) have added further credibility to this theory. Despite the success of the AEM, it is pertinent to remember that it is fundamentally a statistical theory, and other hypotheses (e.g  davidson2006logarithmic; davidson2009simple) can also be used to explain the statistical characteristics of the log layer.
The attached eddy hypothesis is based on the principle that the physics and statistical properties of the logarithmic layer can be explained by considering geometrically self-similar eddies that extend from the wall. A foundational assumption is that the length scale of each individual eddy follows a probability distribution that is a function of the distance from the wall. Hence, the term ‘attached’ alludes to the fact that every eddy can be assumed to be randomly placed on the wall. The AEM is effectively an inviscid theory, yet the range of scales is set by the Reynolds number.
A key advance was made by woodcock2015statistical (henceforth W&M) who established a statistical foundation for AEM. They provide a complete derivation for all the velocity moments and demonstrate logarithmic scaling relationships therein. They were also able to provide expressions for the skewness and flatness of the wall-normal and spanwise fluctuations as a function of the Reynolds number. While variants of the AEM continue to be developed in the literature (e.g. hwang2020attached), we consider W&M as the starting point of our exploration and exclusively consider zero pressure gradient boundary layers. The main contributions and guiding questions in this work are:
  1. 1.
    Inverse identification of attached-eddy influence functions. Townsend introduced the notion of an “eddy intensity” (or contribution) function: a wall-parallel average of the single-eddy induced velocity and its products. While Townsend gave remarkably insightful descriptions of the nature of this function (Figure 5.7, page 155 of townsend1976structure), his and most other treatments are qualitative. Here we pose a concrete inverse problem that infers the ideal influence functions $I_{1}(y/h)$ and $I_{ij}(y/h)$ implied by reference one-point moments (DNS/experiment). The inferred kernels expose the wall-normal support, sign structure, and plateau/decay behavior required for log-layer mean and Reynolds stresses, and motivate a simplified kernel model that yields a direct relation between the von Kármán constant $\kappa$ and the population density $\beta$.
  2. 2.
    A minimal Biot–Savart-consistent attached eddy. W&M utilize an eddy that has a complex shape (Figure 1 in W&M), presumably configured using insight from DNS and/or PIV fields. Guided by the inferred kernels, we construct an attached-eddy template from a small number of Rankine vortex rods (a “square hairpin”) together with an inviscid image system. The objective is not an exact Navier–Stokes solution, but a kinematically consistent building block that (i) respects wall-impermeability, (ii) has well-posed planar influence integrals, and (iii) admits a transparent connection between morphology and the influence functions.
  3. 3.
    What aspects of morphology control log-layer statistics? We quantify the sensitivity of $I_{1}$, $I_{ij}$ and the resulting moments to eddy geometry (e.g. triangular vs. square hairpins, inclination angle, and simple packet/nesting configurations treated as composite marks). This clarifies which structural features are essential (e.g. realizing the near-constant portion of $I_{1}$ for $y/h\lesssim 1$) and which details have weak impact once similarity and attachedness are enforced.
  4. 4.
    A spectral Influence kernel and interpretation of attached-eddy spectra. We introduce an explicit spectral Influence kernel $I_{\phi}(\kappa_{x},\eta)$ that maps a self-similar wall-parallel footprint to its one-dimensional streamwise energy spectrum as a function of relative height $\eta=y/h$. Combining $I_{\phi}$ with the scale-invariant population law $p(h)\propto h^{-3}$ yields a compact explanation for the emergence (and limitations) of the $k_{x}^{-1}$ range, the location of the premultiplied hump, and the origin of low-$k_{x}$ discrepancies under the independent-mark (Poisson) hypothesis..
All forward and inverse modeling tools used here are open sourced [^1] https://github.com/CaslabUM/AttachedEddy.
1  Statistics of attached eddies
We begin with a short description of the attached eddy hypothesis and modeling. This presentation generally follows W&M with a slightly different pedagogy. As noted above, we adopt the standard coordinate system: $x$ streamwise, $y$ wall-normal, and $z$ spanwise. Planar homogenization/influence integrals are taken over the wall-parallel $(x,z)$ plane, so one-point statistics depend only on $y$.
Consider a collection of $n$ eddies (Figure 1) of length scale $h_{e,i}$ that are placed on the wall at locations $\mathbf{x}_{e,i}=(x_{e,i},z_{e,i})$. Since these are attached eddies, it is implicit that only the streamwise and spanwise components of $\mathbf{x}_{e,i}$ are variable. Define $\mathbf{h}_{e}\triangleq\{h_{e,1},h_{e,2},..,h_{e,n}\}$ and $\mathbf{X}\triangleq\{\mathbf{x}_{e,1},\mathbf{x}_{e,2},..,\mathbf{x}_{e,n}\}$. Assuming self-similarity and linearity, a quantity $q$ (e.g. a velocity component) evaluated at a location $\mathbf{x}$ can be determined using the superposition
$q(\mathbf{x},\mathbf{X}_{e},\mathbf{h}_{e},n)\triangleq\sum_{i=1}^{n}q\!\left(\frac{\mathbf{x}-\mathbf{x}_{e,i}}{h_{e,i}}\right).$
Figure 1: Schematic of discrete representation of attached eddies with $n=21$ and $m=3$.
Restricting our attention to a streamwise–spanwise square wall patch of side $2L$ in which the eddies are assumed to be independently and uniformly distributed, and assuming that the range of eddy length scales follows a probability density function $p(h)$, the expectation of $q$ in this region is given by
$q(\mathbf{x},n)\triangleq\mathbb{E}_{\mathbf{X}_{e},\mathbf{h}_{e}}\!\left[q(\mathbf{x},\mathbf{X}_{e},\mathbf{h}_{e},n)\right]=\frac{1}{4L^{2}}\int_{h_{\min}}^{h_{\max}}\int_{-L}^{L}\int_{-L}^{L}\sum_{i=1}^{n}q\!\left(\frac{\mathbf{x}-\mathbf{x}_{e,i}}{h}\right)\,p(h)\,dx_{e,i}\,dz_{e,i}\,dh.$ (1)
Homogenizing in the $(x,z)$ directions, and assuming $L$ is large enough that each eddy centered at the origin has a negligible induced contribution outside an area $4L^{2}$ (Refer Campbell’s theorem rice1944mathematical and the appendix of W&M), it can be shown that
$q(y,n)\triangleq\mathbb{E}_{x,z}\!\left[q(\mathbf{x},n)\right]\approx\frac{n}{4L^{2}}\int_{h_{\min}}^{h_{\max}}\int_{-L}^{L}\int_{-L}^{L}q\!\left(\frac{\mathbf{x}}{h}\right)\,p(h)\,dx\,dz\,dh.$
Note that in the above equation, the entire field is written as a function of one prototypical eddy, scaled by the probability density function of the eddy sizes $p(h)$ and the eddy density $n/(4L^{2})$. If the mean eddy density is $\beta$, then using Poisson’s law, the expected value (over all numbers of eddies) is
$Q(y)\triangleq\mathbb{E}_{n}\!\left[q(y,n)\right]=\beta\int_{h_{\min}}^{h_{\max}}\int_{-L}^{L}\int_{-L}^{L}q\!\left(\frac{\mathbf{x}}{h}\right)\,p(h)\,dx\,dz\,dh.$ (2)
Note that all the $q$’s defined above are random variables, yet $Q(y)$ is a deterministic quantity. Now we are in a position to define the mean streamwise velocity $U(y)$ as a superposition of eddies of various sizes $h$. This can be written in terms of the induced velocity field $u_{1}(\cdot)$ of one prototypical eddy:
$U(y)$ $\triangleq\beta\int_{h_{\min}}^{h_{\max}}\int_{-L}^{L}\int_{-L}^{L}u_{1}\!\left(\frac{\mathbf{x}}{h}\right)\,p(h)\,dx\,dz\,dh+U_{ref}$
$=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\left[\int_{-L/h}^{L/h}\int_{-L/h}^{L/h}u_{1}\!\left(\frac{\mathbf{x}}{h}\right)\,d\!\left(\frac{x}{h}\right)\,d\!\left(\frac{z}{h}\right)\right]dh+U_{ref}$
$=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\,I_{1}\!\left(\frac{y}{h}\right)\,dh+U_{ref},$ (3)
where the mean-flow eddy contribution (influence) function is
$I_{1}\!\left(\frac{y}{h}\right)$ $\triangleq\int_{-L/h}^{L/h}\int_{-L/h}^{L/h}u_{1}\!\left(\frac{\mathbf{x}}{h}\right)\,d\!\left(\frac{x}{h}\right)\,d\!\left(\frac{z}{h}\right).$ (4)
Similarly, we can define the Reynolds stress tensor as
$R_{ij}(y)$ $\triangleq\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\,I_{ij}\!\left(\frac{y}{h}\right)\,dh,$
$\textrm{where }\;I_{ij}\!\left(\frac{y}{h}\right)$ $\triangleq\int_{-L/h}^{L/h}\int_{-L/h}^{L/h}u_{i}\!\left(\frac{\mathbf{x}}{h}\right)\,u_{j}\!\left(\frac{\mathbf{x}}{h}\right)\,d\!\left(\frac{x}{h}\right)\,d\!\left(\frac{z}{h}\right).$
Note the presence of an additional freestream velocity in the definition of $U(y)$. This is required because we are working with induced velocity fluctuations.
The final piece we need is the probability distribution of the eddy sizes. Using insight from townsend1976structure and perry1982mechanism, W&M propose that $p(h)\propto 1/h^{3}$. In particular, for $h\in[h_{\min},h_{\max}]$,
$p(h)\triangleq\frac{C}{h^{3}},\qquad C\triangleq\frac{2}{1/h_{\min}^{2}-1/h_{\max}^{2}}.$
Here $h_{\max}\approx\delta^{+}$ is the outer length scale (boundary layer thickness in wall units) and $h_{\min}$ is set by the friction Reynolds number $Re_{\tau}$. For notational simplicity, all length and velocity scales should be assumed to be in wall units henceforth.
1.1  Statistical assumptions and derivation of the Reynolds-stress formula
The attached-eddy framework above is a shot-noise model: the instantaneous velocity perturbation at a point is written as a superposition of contributions from a random collection of eddies. While the mean follows immediately from linearity of expectation, the Reynolds stresses are quadratic and require an explicit statement of the underlying point-process assumptions.
We idealize the set of attached eddies as a stationary marked Poisson point process $\Phi$ on the wall-parallel plane $(x,z)$. Each eddy is characterized by a wall-parallel center location $\mathbf{x}_{e}=(x_{e},z_{e})$ and a mark $\theta$ that includes its size $h$ (and possibly additional parameters, e.g. orientation, circulation distribution, etc.). The following are assumed:
  1. A1.
    Homogeneity: the process is stationary in $(x,z)$ with constant areal intensity $\beta$ (so the expected number of eddies in area $A$ is $\beta A$).
  2. A2.
    Independent marking: conditional on the locations, the marks $\{\theta_{k}\}$ are i.i.d. and independent of the point locations; in particular the size distribution is $p(h)$.
  3. A3.
    Moment existence: the required single-eddy planar integrals defining $I_{i}$ and $I_{ij}$ exist in the $L/h\to\infty$ limit (see Appendix A for a convergence discussion).
  4. A4.
    Ergodic/homogenization interpretation: planar averaging over a sufficiently large patch is identified with the ensemble mean for the stationary process.
Assumptions A1–A2 are the essential statistical hypotheses behind the Reynolds-stress closure.
Velocity superposition.
Let a single prototypical eddy with mark $\theta$ induce a velocity field $u_{i}^{(e)}(\xi,\zeta,\eta;\theta)\,$ in dimensionless coordinates $(\xi,\zeta,\eta)=(x/h,\;z/h,\;y/h).$ Define the single-eddy planar influence integral
$I_{i}(\eta;\theta)\;\triangleq\;\int_{\mathbb{R}^{2}}u_{i}^{(e)}(\xi,\zeta,\eta;\theta)\,d\xi\,d\zeta,\qquad(\eta=y/h).$ (5)
Under A1–A3, first-order Campbell/Mecke yields
$U_{i}(y)\;\triangleq\;\mathbb{E}[u_{i}(y)]\;=\;\beta\int p(h)\,h^{2}\,\mathbb{E}_{\theta|h}\!\left[I_{i}\!\left(\frac{y}{h};\theta\right)\right]\,dh.$ (6)
Second moment and the cross-term issue.
The central subtlety is that $u_{i}u_{j}=(\sum_{k}u_{i}^{(k)})(\sum_{\ell}u_{j}^{(\ell)})$ contains cross terms with $k\neq\ell$, so one cannot in general write the second moment as a sum of “individual Reynolds stresses”. To make this explicit, expand
$\mathbb{E}[u_{i}(y)u_{j}(y)]=\mathbb{E}\!\left[\sum_{k}u_{i}^{(k)}u_{j}^{(k)}\right]+\mathbb{E}\!\left[\sum_{k\neq\ell}u_{i}^{(k)}u_{j}^{(\ell)}\right].$ (7)
Define the single-eddy quadratic kernel
$I_{ij}(\eta;\theta)\;\triangleq\;\int_{\mathbb{R}^{2}}u_{i}^{(e)}(\xi,\zeta,\eta;\theta)\,u_{j}^{(e)}(\xi,\zeta,\eta;\theta)\,d\xi\,d\zeta.$ (8)
For a Poisson process with independent marks (A1–A2), the second-order Campbell theorem gives the closed form
$\mathbb{E}[u_{i}(y)u_{j}(y)]=\beta\int p(h)\,h^{2}\,\mathbb{E}_{\theta|h}\!\left[I_{ij}\!\left(\frac{y}{h};\theta\right)\right]\,dh\;+\;U_{i}(y)\,U_{j}(y),$ (9)
where the $U_{i}U_{j}$ term is precisely the contribution of the $k\neq\ell$ cross terms in (7). In other words, under A1–A2, the cross terms do not vanish individually; rather, they factorize into the product of means.
Reynolds stresses as covariances (cancellation of the cross term).
The one-point Reynolds stress tensor is the covariance of the total field:
$R_{ij}(y)\;\triangleq\;\mathbb{E}\!\left[(u_{i}(y)-U_{i}(y))(u_{j}(y)-U_{j}(y))\right]=\mathbb{E}[u_{i}(y)u_{j}(y)]-U_{i}(y)U_{j}(y).$ (10)
Substituting (9) into (10) yields the single-eddy representation
$R_{ij}(y)=\beta\int p(h)\,h^{2}\,\mathbb{E}_{\theta|h}\!\left[I_{ij}\!\left(\frac{y}{h};\theta\right)\right]\,dh.$ (11)
Equation (11) is therefore justified only when the eddy field is modeled as a Poisson (or otherwise “factorizable”) superposition with independent marks.
If the eddy process is not Poisson (e.g. clustering, exclusion, or explicit packet/nesting correlations), then the cross term in (7) no longer factorizes into $U_{i}U_{j}$. In that case, (9) acquires an additional correction involving the pair (second factorial) moment measure of the process (equivalently a pair-correlation function), and the covariance is no longer determined solely by the single-eddy kernel $I_{ij}$. A statistically consistent way to incorporate a correlated internal structure (e.g. a hairpin packet) while retaining (11) is to treat the packet as the Poisson object: internal correlations are absorbed into the mark $\theta$, while packet centers remain a stationary Poisson field.
1.2  Convergence of planar influence integrals
Equation 4 defines the mean-flow eddy contribution (influence) function as a planar integral of the streamwise velocity induced by a single prototypical eddy. Writing the dimensionless coordinates $(\xi,\zeta,\eta)\triangleq(x/h,\;z/h,\;y/h)$ and $R\triangleq L/h$, we can make the dependence on the truncation radius explicit:
$I_{1}(\eta;R)\;\triangleq\;\int_{-R}^{R}\int_{-R}^{R}u_{1}(\xi,\zeta,\eta)\,d\xi\,d\zeta,\qquad I_{1}(\eta)\;\triangleq\;\lim_{R\to\infty}I_{1}(\eta;R),$ (12)
whenever the limit exists. In particular, we do not assume compact support of the induced velocity field. Rather, the required condition for Eq. 4 to be well-defined is that the planar integral converges as $R\to\infty$, or equivalently that the tail contribution beyond $|(\xi,\zeta)|\gtrsim R$ becomes negligible for the chosen truncation.
Let $\rho\triangleq\sqrt{\xi^{2}+\zeta^{2}}$. A standard sufficient condition for convergence is that
$\int_{\mathbb{R}^{2}}|u_{1}(\xi,\zeta,\eta)|\,d\xi\,d\zeta<\infty.$ (13)
A convenient and interpretable sufficient condition is a power-law far-field bound $|u_{1}(\xi,\zeta,\eta)|\leq C(\eta)\,(1+\rho)^{-p}$ for some $p>2$. Indeed, the tail can be bounded (up to an angular factor) by
$\big|I_{1}(\eta)-I_{1}(\eta;R)\big|\;\leq\;\int_{\rho>R}|u_{1}|\,dA\;\lesssim\;2\pi C(\eta)\int_{R}^{\infty}\rho^{1-p}\,d\rho\;=\;\frac{2\pi C(\eta)}{p-2}\,R^{2-p}.$ (14)
Hence, $p>2$ guarantees convergence as $R\to\infty$. The borderline case $p=2$ produces a logarithmic divergence $\sim\log R$; for $p=3$ the tail decays as $\mathcal{O}(1/R)$.
Reynolds-stress kernels are less restrictive.
Similarly, the second-moment influence functions are
$I_{ij}(\eta;R)\;\triangleq\;\int_{-R}^{R}\int_{-R}^{R}u_{i}(\xi,\zeta,\eta)\,u_{j}(\xi,\zeta,\eta)\,d\xi\,d\zeta,\qquad I_{ij}(\eta)\;\triangleq\;\lim_{R\to\infty}I_{ij}(\eta;R),$ (15)
whenever the limit exists. A sufficient condition for $I_{ij}$ to be well-defined is $u_{i}(\cdot,\cdot,\eta)\,u_{j}(\cdot,\cdot,\eta)\in L^{1}(\mathbb{R}^{2})$. If each velocity component satisfies the far-field decay $|u_{k}|\lesssim\rho^{-p}$, then $u_{i}u_{j}\lesssim\rho^{-2p}$, and the planar tail behaves like $\int_{R}^{\infty}\rho\,\rho^{-2p}\,d\rho=\int_{R}^{\infty}\rho^{1-2p}\,d\rho$, which converges whenever $p>1$. Thus, the convergence requirement for the mean-flow influence $I_{1}$ (which requires $p>2$) is typically the more restrictive requirement; once $I_{1}$ is well-defined, the second-moment kernels $I_{ij}$ are generally well-defined for the same eddy field.
Implications for Biot–Savart line-vortex eddies and the role of the image system.
For line-vortex Biot–Savart fields, a generic open filament can generate a $\rho^{-2}$ contribution to certain velocity components in the far field, which is precisely the borderline case for $I_{1}$ and would lead to a logarithmic growth of $I_{1}(\eta;R)$ with $R$. The specific square-hairpin eddy used herein is paired with an opposite-sign mirror image across $y=0$ (to enforce inviscid wall boundary conditions), and the wall-parallel segment cancels with its image. This construction removes leading-order contributions responsible for a $\rho^{-2}$ tail in the streamwise component and yields a faster far-field decay. In other words, the image construction is not a technical convenience: it is central to ensuring that Eq. 4 is well posed for a line-vortex eddy model.
2  The ideal eddy contribution function
The functions $I_{1}(y/h)$ and $I_{ij}(y/h)$ introduced in Eq. (4) are key for the moments produced by the attached-eddy model: once these single-eddy planar integrals are specified (together with $\beta$ and $p(h)$), the attached-eddy integrals yield the one-point mean and Reynolds stresses, and (through similar constructions) higher-order moments. Rather than prescribing a particular eddy geometry a priori, we first ask a more direct question:
Given one-point statistics $U_{\rm data}(y)$ and $R_{ij,{\rm data}}(y)$ (from DNS or experiment), what shape of $I_{1}(\eta)$ and $I_{ij}(\eta)$, with $\eta=y/h$, is implied by the attached-eddy integral relations?
This is an inverse problem for a Fredholm integral equation of the first kind. In its continuous form,
$U(y)-U_{ref}$ $=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\,I_{1}\!\left(\frac{y}{h}\right)\,dh,$ (16)
$R_{ij}(y)$ $=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,h^{2}\,I_{ij}\!\left(\frac{y}{h}\right)\,dh.,$ (17)
where $U_{ref}$ is the outer velocity (e.g. center line or freestream velocity). For the classical attached-eddy size distribution $p(h)=C/h^{3}$, these become
$U(y)-U_{ref}=\beta\,C\int_{h_{\min}}^{h_{\max}}\frac{I_{1}(y/h)}{h}\,dh,\qquad R_{ij}(y)=\beta\,C\int_{h_{\min}}^{h_{\max}}\frac{I_{ij}(y/h)}{h}\,dh,$ (18)
and, after the change of variables $\eta=y/h$ (so that $dh/h=-d\eta/\eta$),
$U(y)-U_{ref}=\beta\,C\int_{y/h_{\max}}^{y/h_{\min}}\frac{I_{1}(\eta)}{\eta}\,d\eta,\qquad R_{ij}(y)=\beta\,C\int_{y/h_{\max}}^{y/h_{\min}}\frac{I_{ij}(\eta)}{\eta}\,d\eta.$ (19)
Equation (19) makes two important points transparent: (i) the mapping $I(\eta)\mapsto U(y)$ is smoothing (ill-conditioned to invert without regularization), and (ii) a plateau of $I_{1}(\eta)$ over $\eta<1$ immediately yields logarithmic dependence of $U(y)$ on $y$.
2.1  Discrete formulation and identifiability
In practice, we discretize the eddy-size range using $\tilde{\mathbf{h}}=\{h_{1},\ldots,h_{m}\}$ (log-spaced between $h_{\min}$ and $h_{\max}$) and sample the wall-normal coordinate at $\tilde{\mathbf{y}}=\{y_{1},\ldots,y_{m_{y}}\}$. The forward map for the mean can be written compactly as
$\mathbf{U}(\tilde{\mathbf{y}})-U_{ref}\mathbf{1}=\beta\,\mathbf{K}(\mathbf{c})\mathbf{b},$ (20)
where $\mathbf{c}$ contains the unknown nodal values of $I_{1}(\eta)$ (in a chosen basis) and $\mathbf{K}$ is the corresponding discretized integral operator (assembled from the weights $p(h)\,h^{2}$ and the ratios $\eta=y/h$). An analogous system holds for each component $I_{ij}$.
The discretized velocity moments are
$U(y_{r};\mathbf{d})$ $=\beta\sum_{l=1}^{m}p(h_{l})h_{l}^{2}\left[\sum_{q}\sum_{p}u_{1}\left(\frac{{x_{p},z_{q},y_{r}}}{h_{l}};\mathbf{d}\right)\left(\frac{\Delta x}{h_{l}}\right)\left(\frac{\Delta z}{h_{l}}\right)\right]\Delta h+U_{ref}$
$R_{ij}(y_{r};\mathbf{d})$ $=\beta\sum_{l=1}^{m}p(h_{l})h_{l}^{2}\left[\sum_{q}\sum_{p}u_{i}\left(\frac{{x_{p},z_{q},y_{r}}}{h_{l}};\mathbf{d}\right)u_{j}\left(\frac{{x_{p},z_{q},y_{r}}}{h_{l}};\mathbf{d}\right)\left(\frac{\Delta x}{h_{l}}\right)\left(\frac{\Delta z}{h_{l}}\right)\right]\Delta h.$
Explicitly $K_{ij}=I_{1}(z_{i}/h_{j})$ and $b_{i}=\beta p(h_{i})h_{i}^{2}\Delta h.$ We would now like to extract these coefficients. Towards this end, we define the eddy contribution function at selected locations, and interpolate for the values in other locations in a piecewise linear fashion. In other words
$\mathbf{K}=\begin{bmatrix}I_{1}(1)&I_{1}(h_{1}/h_{2})&I_{1}(h_{1}/h_{3})&..&I_{1}(h_{1}/h_{m-1})&I_{1}(h_{1}/h_{m})\\ I_{1}(h_{2}/h_{1})&I_{1}(1)&I_{1}(h_{2}/h_{3})&..&I_{1}(h_{2}/h_{m-1})&I_{1}(h_{2}/h_{m})\\ ..&..&..&..&..&..\\ I_{1}(h_{m}/h_{1})&I_{1}(h_{m}/h_{2})&I_{1}(h_{m}/h_{3})&..&I_{1}(h_{m}/h_{m-1})&I_{1}(1)\end{bmatrix}$
$\triangleq\begin{bmatrix}c_{1}&c_{m+1}&c_{m+2}&..&c_{2m-2}&c_{2m-1}\\ c_{2}&c_{1}&I_{1}(h_{2}/h_{3})&..&I_{1}(h_{2}/h_{m-1})&I_{1}(h_{2}/h_{m})\\ ..&..&..&..&..&..\\ c_{m}&I_{1}(h_{m}/h_{2})&I_{1}(h_{m}/h_{3})&..&I_{1}(h_{m}/h_{m-1})&c_{1}\\ \end{bmatrix}$.
The unknown $I_{1}(\cdot)$ values are then interpolated  [^2] For instance, if $h_{\min}=186$ and $h_{\max}=5186$ and $m=11$, then $I_{1}(h_{2}/h_{3})=0.42106c_{1}+0.57894c_{12}$ from the nodal locations $\mathbf{c}$. Note that there are more unknowns ($2m$) than equations ($m$), and so while it is possible to determine the unknowns that lead to perfect match of a reference (DNS or experiment) velocity profiles, a choice has to be made on the problem formulation. We choose to solve the following minimum norm least squares problem:
$\mathbf{c}_{\textrm{opt}}=\min\mathbf{c}^{T}\mathbf{c}\ \ \textrm{such that}\ \ \mathbf{U}_{ref}(\tilde{\mathbf{y}})-U_{ref}=\mathbf{K}\mathbf{b}.$
Figure 2 shows the optimal eddy contribution functions inferred from the channel flow data (lee2015direct) at friction Reynolds number $Re_{\tau}\approx 5200$. Figure 3 confirms that the first and second velocity moments are perfectly reproduced across the entire channel.
Figure 2: Optimal influence functions for $Re_{\tau}\approx 5200$ for the mean flow (left) and Reynolds stresses (right, with red=streamwise; green=spanwise; blue=wall-normal, and black=shear).
Figure 3: Reference (symbols) vs optimal attached eddy statistics for $Re_{\tau}\approx 5200$
2.2  A simplified influence-function model and the log law
We remark that the underlying physics is more complex than attached eddies can represent. The inverse problem, however, yields a function that, while smooth, has a visually interpretable structure: a near-plateau over $\eta=y/h\lesssim 1$ followed by a transition and decay for $\eta>1$. This motivates the simple piecewise model shown in Fig. 4.
Consider a collection of attached eddies that result in a simple function as shown in Fig. 4.
Figure 4: A hypothetical model of the eddy influence function corresponding to the mean streamwise velocity (blue dashed lines) compared to the optimal influence function.
It will be shown below that it is indeed possible to construct an attached eddy that corresponds to such a function. Using this influence function, one can reconstruct the mean streamwise velocity as
$U(y)$ $=\frac{2\beta}{1/h_{\min}^{2}-1/h_{\max}^{2}}\int_{h_{min}}^{h_{max}}\frac{I_{1}\left({y}/{h}\right)}{h}dh+U_{ref}$
$=\frac{2\beta}{1/h_{\min}^{2}-1/h_{\max}^{2}}\int_{y/h_{max}}^{y/h_{min}}\frac{I_{1}\left({y}/{h}\right)}{y/h}d(y/h)+U_{ref}$
$\approx\frac{2\beta}{1/h_{\min}^{2}-1/h_{\max}^{2}}\left[\int_{y/h_{max}}^{1}\frac{-a_{0}}{y/h}d(y/h)+\int_{1}^{a_{1}/a_{2}}\frac{a_{1}-a_{2}y/h}{y/h}d(y/h)\right]+U_{ref}$
$=\frac{2\beta}{1/h_{\min}^{2}-1/h_{\max}^{2}}\left[a_{0}\log[y/h_{\max}]+a_{1}\log[a_{1}/a_{2}]-a_{1}+a_{2}\right]+U_{ref}.$
It is thus clear that the Kármán constant can be constructed as
$\kappa=\frac{1/h_{\min}^{2}-1/h_{\max}^{2}}{2a_{0}\beta}.$ (21)
Note that W&M use a Taylor series approximation on a generic eddy to derive an alternate expression for $\kappa$. While that is insightful, it is perhaps a valid approximation only for certain types of eddy influence functions.
3  Predictions
Now, we answer the question whether it is possible to recreate the above hypothetical influence function using an attached eddy. Consider a hairpin (Figure 5) with unit circulation as the attached eddy, along with its image across the $y=0$ plane to enforce no-penetration. We consistently achieved good results using Rankine vortex-rods instead of vortex filaments. All the results henceforth correspond to vortex segments represented as Rankine vortex rods with core radius $\approx 0.025$ units relative to Fig. 5.
It is remarked that this eddy corresponds to a closed loop in the sense that the leg of the eddy aligned with the wall is canceled by an equal and opposite image vortex pair, and is thus consistent with the vorticity kinematics. Using the Biot–Savart law to compute the induced velocity $u_{1}$ in Eqn. 4, Figure 6 shows that such a simple attached eddy can reproduce the mean streamwise velocity accurately. Further, the second moments are also seen to be well-predicted as shown in Figure 6.
Figure 5: A prototypical hairpin-type eddy.
Figure 6: Velocity moments and eddy influence function. Lines: Optimal eddy influence function. Symbols: Square hairpin at $\theta=60^{\circ}$. For the right figures, colors follow Figure 2.
It is noted that an extensive optimization was performed on different eddy shapes (e.g. triangular hairpin) and even packet configurations. Yet, these did not yield improved predictions. The hairpin at an angle of $\theta=68^{\circ}$ appears to be most optimal. We now further evaluate the predictive capabilities of this model by using the $\theta=60^{\circ}$ configuration and comparing it to the experimental measurements from samie2018fully. As shown in Figures 7,8,9, the model remains remarkably accurate in predicting mean velocity, streamwise Reynolds stress and spectra over a range of Reynolds numbers.
Figure 7: Mean velocity prediction using square hairpin at $\theta=60^{\circ}$ for $Re_{\tau}=[6000,10000,14500,20000]$. Symbols :  samie2018fully
Figure 8: Streamwise velocity fluctuation predictions using square hairpin at $\theta=60^{\circ}$ for $Re_{\tau}=[6000,10000,14500,20000]$. Symbols :  samie2018fully
Figure 9: Pre-multiplied streamwise energy spectra predictions at $y^{+}=1184$ using square hairpin at $\theta=60^{\circ}$ for $Re_{\tau}=[6000,10000,14500,20000]$. Symbols :  samie2018fully
4  Physical and Statistical Insights
Once the eddy shape is chosen and $\beta$ is fixed, there are no degrees of freedom left. Yet, the comparisons in the previous section suggest that a minimal hairpin-type eddy can reproduce not only the one-point moments at $Re_{\tau}\approx 5200$ but also the evolution of $U^{+}$, $u^{\prime 2}$ and the premultiplied spectrum across Reynolds numbers. This section provides a more substantial interpretation of why this occurs, which ingredients appear essential, and which limitations are inherent to an independent-eddy superposition.
It is useful to separate (i) the population law $(\beta,p(h),h_{\min},h_{\max})$ from (ii) the single-eddy template (geometry, inclination angle $\theta$, circulation scaling, and vortex-core regularization). In this study we adopt Townsend/Perry scaling $p(h)\propto h^{-3}$ and determine $\beta$ from the log-law relation (Eq. 21) using $\kappa=0.384$, after which $\beta$ is held fixed across the Reynolds-number comparisons in Section 3. The template is likewise fixed once (square hairpin with $\theta=60^{\circ}$ and a small Rankine core), so the subsequent moment and spectral comparisons should be interpreted as out-of-sample tests of that single-eddy choice rather than repeated parameter fitting. Under these conventions, the only continuously adjusted parameter used to anchor the data is the population amplitude $\beta$.
The inverse problem strongly constrains what any viable attached eddy must do in a wall-parallel plane. In particular, the inferred mean influence function $I_{1}(y/h)$ exhibits (a) a near-constant region for $y/h\lesssim 1$ and (b) rapid decay for $y/h\gtrsim 1$, implying that eddies larger than the observation height dominate the log-region population integral. The square hairpin + image realizes these features because the two legs and head generate partially canceling far-field contributions, while the inviscid image enforces wall compatibility and suppresses the leading-order $1/\rho^{2}$ tail that would otherwise make the planar influence integrals marginal (Appendix A). By contrast, a triangular hairpin produces a more monotone “ramp” in $I_{1}$ and does not reproduce the inferred plateau as cleanly (Figure 6). Experimentation with a pack of triangular hairpins with variable circulation strengths did not improve results beyond the square hairpin which appears optimal.
The inclined hairpin footprint produces elongated streamwise $u$ signatures, which is precisely the kind of wall-parallel signature that feeds the low-$k_{x}$ premultiplied hump. The log-layer statistics depend weakly on microscopic details once a few structural constraints are met. It is possible that this eddy structure has some key features that place it in a robust universality class for log-layer statistics. We examine this below by providing an explicit similarity map from eddy morphology to the single-eddy spectrum as a function of relative height.
4.1  Similarity structure and the Influence kernel
The planar influence functions $I_{1}(y/h)$ and $I_{ij}(y/h)$ govern one-point moments via Eq. (3) and Eq. (11). For spectra, we introduce an analogous spectral Influence kernel $I_{\phi}(\kappa_{x},\eta)$ that compactly summarizes how a single self-similar eddy distributes streamwise energy across wavenumbers as a function of relative height $\eta=y/h$. To the author’s knowledge, making this kernel explicit is not common in the attached-eddy literature; it is useful because it cleanly separates single-eddy geometry (captured by $I_{\phi}$) from population effects (captured by $p(h)$ and $\beta$).
Under our geometric and circulation scaling assumptions, the hairpin footprint is self-similar:
$u_{h}(x,z;y)=\mathcal{U}\!\left(\frac{x}{h},\frac{z}{h},\frac{y}{h}\right)=\mathcal{U}(\xi,\zeta,\eta),$
where $\mathcal{U}$ is the unit-eddy footprint expressed in dimensionless coordinates. Taking the Fourier transform and changing variables yields the exact scaling $\widehat{u}_{h}(k_{x},k_{z};y)=\,h^{2}\,\widehat{\mathcal{U}}(\kappa_{x},\kappa_{z};\eta),$ where $\widehat{\mathcal{U}}(\kappa_{x},\kappa_{z};\eta)=\int_{\mathbb{R}^{2}}\mathcal{U}(\xi,\zeta,\eta)\,e^{-i(\kappa_{x}\xi+\kappa_{z}\zeta)}\,d\xi\,d\zeta.$ Therefore
$\Phi_{xx}^{(h)}(k_{x},k_{z};y)=\frac{\,h^{4}}{(2\pi)^{2}}\,\left|\widehat{\mathcal{U}}(\kappa_{x},\kappa_{z};\eta)\right|^{2}.$
Integrating over $k_{z}$ and using $k_{z}=\kappa_{z}/h$ gives the scalewise spectrum
$\phi_{xx}^{(h)}(k_{x};y)\triangleq\int_{-\infty}^{\infty}\Phi_{xx}^{(h)}(k_{x},k_{z};y)\,dk_{z}=\,h^{3}\,I_{\phi}(\kappa_{x},\eta),$
where the Influence kernel is defined by
$I_{\phi}(\kappa_{x},\eta)\triangleq\frac{1}{(2\pi)^{2}}\int_{-\infty}^{\infty}\left|\widehat{\mathcal{U}}(\kappa_{x},\kappa_{z};\eta)\right|^{2}\,d\kappa_{z}.$
This identity is a concrete mathematical object that will help decode the energy spectrum in greater detail.
Figure 10: Streamwise energy Influence kernel $I_{\phi}(\kappa_{x},\eta)$ for square hairpin at $\theta=60^{\circ}$. This figure highlights that the kernel has a broad, elevated region at small $\eta$ and small-to-moderate $\kappa$, and a more localized hump as $\eta$ increases. Also shown are contributions to $\phi_{xx}^{+}(k_{x}^{+})$ for $k_{x}^{+}=1\times 10^{-4}$ (magenta) and $k_{x}^{+}=1\times 10^{-3}$ (red) at $y^{+}=1184$ for $Re_{\tau}=20000$.
The kernel $I_{\phi}(\kappa_{x},\eta)$ is a function of the dimensionless wavenumber $\kappa_{x}=k_{x}h$ at fixed relative height $\eta=y/h$ and is shown in Figure 10. It is immediately clear that eddies of scale $h<y^{+}$ have negligible influence on the statistics.
For any fixed $\eta$, typical scaling arguments yield a concentration the influence $S(\kappa_{x},\eta)$ around $\kappa_{x}=O(1)$ because $\kappa_{x}$ is the only dimensionless streamwise wavenumber available to the unit eddy. Therefore, for a given size $h$ and fixed $y$, the dominant contribution of that eddy appears around
$k_{x}h=O(1)\quad\Longleftrightarrow\quad k_{x}=O(1/h),$
equivalently $\lambda_{x}=2\pi/k_{x}=O(h)$. This explains the peak of the premultiplied spectra in Fig. 9. Additionally, the square hairpin Attached Eddy model is able to recover a near-perfect linear behavior of the streamwise energy spectrum in the intermediate scale range as shown in Fig. 11. Below, we will give more concrete and quantitative explanations.
Figure 11: Streamwise energy spectra at $y^{+}=1184$ for $Re_{\tau}=20000$.
4.2  From single-eddy spectra to the attached-eddy spectrum
The attached-eddy prediction for the 1D spectrum is the population superposition
$\phi_{xx}(k_{x};y)=\beta\int_{h_{\min}}^{h_{\max}}p(h)\,\phi_{xx}^{(h)}(k_{x};y)\,dh,\qquad p(h)=\frac{C}{h^{3}}.$
Insert the exact single-eddy factorization:
$\phi_{xx}(k_{x};y)=\beta\int_{h_{\min}}^{h_{\max}}\frac{C}{h^{3}}\,\big(\,h^{3}\,I_{\phi}(k_{x}h,y/h)\big)\,dh=\beta\,C\,\int_{h_{\min}}^{h_{\max}}I_{\phi}(k_{x}h,y/h)\,dh.$
Change variables $\kappa=k_{x}h$ so $dh=d\kappa/k_{x}$, and note that $y/h=yk_{x}/\kappa$:
$\phi_{xx}(k_{x};y)=\frac{\beta\,C\,}{k_{x}}\int_{k_{x}h_{\min}}^{k_{x}h_{\max}}I_{\phi}\!\left(\kappa,\frac{yk_{x}}{\kappa}\right)\,d\kappa.$
Define the integral factor
$F(k_{x};y)=\int_{k_{x}h_{\min}}^{k_{x}h_{\max}}I_{\phi}\!\left(\kappa,\frac{yk_{x}}{\kappa}\right)\,d\kappa,$
so that
$\phi_{xx}(k_{x};y)=\frac{\beta\,C\,}{k_{x}}\,F(k_{x};y).$
The $1/k_{x}$ scaling in the logarithmic region arises when $F(k_{x};y)$ varies slowly with $k_{x}$ over an intermediate band. This occurs when the $\kappa$-window $k_{x}h_{\min},k_{x}h_{\max}]$ covers the portion of the kernel $S(\kappa,\eta)$ that carries most of the contribution, and when the effective $\eta$ values encountered (through $\eta=y/h=yk_{x}/\kappa$) lie in the log-layer-relevant range. In that band,
$F(k_{x};y)\approx\text{constant}\quad\Longrightarrow\quad\phi_{xx}(k_{x};y)\propto\frac{1}{k_{x}}\propto\lambda_{x}.$
This is substantiated by the results in Fig. 12.
Figure 12: Contribution to $F$ integral from different scales at $y^{+}=1184$ for $Re_{\tau}=20000$. The smallest-$k_{x}$ trajectory samples the left/low-$\kappa$ edge of the kernel and remains at small $\eta=y/h$.
The above $1/k_{x}$” argument corresponds to the case that $F(k_{x};y)$ is approximately constant over an intermediate $k_{x}$ range. However the low-$k_{x}$ behavior is entirely controlled by how $F(k_{x};y)$ varies as $k_{x}\to 0$. There is no universality guarantee there: when $k_{x}$ is small, the $\kappa$-window and the $\eta$-trajectory can force the integral to sample the least robust part of the kernel. In other words, for small $k_{x}$, this trajectory spends substantial “arc length” in the region of small $\kappa$ and small $\eta$.
This is visible in the kernel visualization. Figure 10 shows the kernel surface, and Figure 12 shows three trajectories corresponding to fixed $k_{x}^{+}$ values. The low-$\kappa$ region of the kernel is where $I_{\phi}(\kappa,\eta)$ typically changes most rapidly. In such a regime, the integral $F(k_{x};y)$ cannot be expected to be constant: small shifts in $k_{x}$ move both the integration bounds and the trajectory through $(\kappa,\eta)$, producing non-negligible changes in $F$.
Physically, low $k_{x}$ corresponds to very long streamwise wavelengths (VLSM-like content). In an attached-eddy framework with independent marks (Poisson placement in the wall-parallel plane), the low-$k_{x}$ variance is built by incoherently summing contributions of the largest eddies. In a Poisson superposition, long-wavelength modes are fed by many large eddies with essentially random phases, so variances add without cancellation. Real wall turbulence exhibits organization, intermittency, and correlations at the largest scales; those effects can reduce energy in the very lowest modes relative to an independent-eddy superposition.  baars2020data1; baars2020data2 offer more quantitative insights in the form of 3 different eddy types, and that attached eddies represent one of these types and contribute a large fraction of the energy only at a intermediate range of scales in the log layer. The latter point has been well argued by  perry1995wall. The independent-mark hypothesis therefore tends to overpopulate the lowest wavenumbers unless correlations are explicitly modeled or packaged into larger marks. It is possible that the influence kernel $I_{\phi}$ can be used in an inverse problem setting to ‘design’ compositions of prototypical eddies to model desired spectral properties.
5  Related approaches and connections
The present work occupies a complementary position within the broader landscape of attached-eddy theories and coherent-structure studies. An important parallel development is the hierarchical random additive process (HRAP) model introduced by YangMarusicMeneveau2016HRAP which provides an alternative statistical reinterpretation of Townsend’s attached eddy hypothesis. This formulation elegantly recovers logarithmic scaling laws for moments, structure functions, and generalized two-point correlations without requiring explicit specification of eddy geometry.
In this sense, the HRAP model and the marusic2019attached review reinforce a key message: the precise morphology of the prototypical eddy is not essential for reproducing log-layer statistics. The present work provides a concrete mechanistic underpinning for this observation. The inverse-problem analysis (Section 2) demonstrates that the inferred influence functions $I_{1}(\eta)$ and $I_{ij}(\eta)$ represent the essential kinematic constraints that any viable eddy must satisfy. Once these constraints are met: a near-plateau for $\eta\lesssim 1$ and appropriate decay for $\eta>1$, the detailed vortical architecture becomes secondary. The square hairpin succeeds not because it is the unique physical structure, but because it is a minimal template that satisfies these influence-function requirements while remaining consistent with Biot–Savart kinematics and wall boundary conditions. The HRAP framework abstracts this further by encoding only the statistical consequence (additive contributions from a hierarchy of scales) without committing to any particular morphology.
It is important to distinguish the present statistical/kinematic AEM from the growing body of work that seeks to identify wall-attached eddies as instantaneous, material features in DNS or experimental data. Beginning with delAlamo2006Clusters who identified self-similar vortex clusters in the logarithmic region of turbulent channels, a series of studies have extracted three-dimensional coherent structures using various detection criteria. DennisNickels2011 provided detailed experimental measurements of vortex packets in turbulent boundary layers via tomographic PIV. LozanoDuranFloresJimenez2012 characterized the three-dimensional structure of momentum-transferring Q2/Q4 events, demonstrating that attached structures become geometrically self-similar with sizes proportional to their distance from the wall. More recently, HwangSung2018 developed a clustering methodology to extract wall-attached structures of velocity fluctuations, and Cheng2019Identity; Cheng2020Uncovering employed bidimensional empirical mode decomposition to isolate attached eddies and quantify their contribution to skin-friction generation.
These data-driven approaches validate the physical premise of wall-attached, self-similar hierarchies, providing essential structures that models such as the AEM seek to represent statistically. However, they also reveal complexity that the minimal AEM intentionally abstracts away: instantaneous eddies exhibit significant variability in shape, orientation, and internal structure; they interact nonlinearly; and their dynamics involve processes (generation, merging, cascade) not captured by easily interpretable models. The present work thus occupies a middle ground: it provides more geometric specificity than purely statistical models (HRAP) while remaining deliberately simpler than the full complexity revealed by structure-extraction studies. The spectral influence kernel $I_{\phi}(\kappa_{x},\eta)$ introduced in Section 4.1 offers a bridge between these perspectives and encodes the spectral signature of a single eddy template and can, in principle, be compared against conditionally averaged spectra from identified structures. Such comparisons, along with extensions to two-dimensional $(k_{x},k_{z})$ spectra and cross-component correlations, represent promising directions for strengthening the connection between kinematic models and data-driven structure identification.
davidson2009simple propose a deliberately simple log-layer model for streamwise fluctuations which reproduces the $k_{x}^{-1}$ behavior of the one-dimensional longitudinal spectrum in the log region and provides reasonable predictions for the variance $u^{\prime}u^{\prime}$ (up to an additive constant). A key conceptual message is that log-layer streamwise statistics are strongly constrained by (i) attachedness and (ii) scale-invariant population arguments, so one-dimensional spectra and even-order statistics can be surprisingly insensitive to the fine morphological details of the underlying structures. In that sense, the Davidson–Krogstad viewpoint supports the observation that once the mean and a small set of scaling assumptions are fixed, the streamwise spectrum in the log region has limited freedom.
6  Conclusions and Perspectives
This work extended the statistical foundations of the Attached eddy modeling framework by clarifying the formulation and offering proofs of convergence. Additional contributions of this work can be summarized as follows:
  • Inverse problem for attached-eddy kernels. We formulate the recovery of the single-eddy influence functions $I_{1}(y/h)$ and $I_{ij}(y/h)$ from reference one-point moments as an explicit inverse problem (Section 2). This makes the “eddy contribution function” a quantitative, inferable object rather than a qualitative sketch.
  • Ideal-kernel structure. For the $Re_{\tau}\approx 5200$ DNS considered, the inferred mean-flow kernel exhibits a near-plateau over $y/h\lesssim 1$ and a transition/decay for $y/h>1$ (Figure 2). This is precisely the structural feature required for logarithmic scaling in Eq. (19).
  • Minimal Biot–Savart-consistent eddy template. A simple square hairpin-type vortex-loop template (implemented with Rankine vortex rods) together with an inviscid image system across the wall can reproduce both the inferred kernels and the corresponding mean and Reynolds-stress profiles with good accuracy (Figure 6).
  • A spectral Influence kernel. We introduce the streamwise-energy Influence kernel $I_{\phi}(\kappa_{x},\eta)$ (Section 4.1), which provides a compact representation of how a self-similar eddy footprint populates the one-dimensional spectrum as a function of scale. The spectral content of a single self-similar eddy as a function of dimensionless wavenumber and relative height is implicit in prior work, possibly dating back to  perry1982mechanism. In this work, the kernel makes the emergence of the $k_{x}^{-1}$ scaling band and its limitations explicit, and more transparent than population-level arguments alone and offers insight into constructing more sophisticated compositional eddy models. It is notable that $\int_{\kappa_{x}}I_{\phi}(\kappa_{x},\eta)d\kappa_{x}=I_{11}(\eta).$
The original goal of this work was to use the inferred influence functions to uncover an optimal eddy or packet configuration. One of the conclusions of this work is that a simple Rankine rod-based square hairpin is unusually predictive once the mean is anchored, and is able to represent attached eddy physics. The reasons are as follows:
  • The single-eddy spectrum is determined by $I_{\phi}(\kappa_{x},\eta)$, a two-variable kernel. Population integration introduces only the scale window and one amplitude constant ($\beta$).
  • $p(h)\propto h^{-3}$ cancels the $h^{3}$ prefactor exactly. This produces the $1/k_{x}$ factor explicitly and shifts the remaining complexity into the slowly varying integral $F(k_{x};y)$.
  • The tilted hairpin footprint produces streamwise-extended $u$ signatures on a wall-parallel plane. Through the $k_{x}h$ mapping, the dominant wavelengths at $y$ are in the decade $\lambda_{x}=O(10y)$, which at $y^{+}\sim 10^{3}$ sits near $\lambda_{x}^{+}\sim 10^{4}$.
  • Reynolds-number variation enters mainly through the window endpoints. For fixed $y^{+}$ in the log region, changing $Re_{\tau}$ changes $h_{\max}$ and therefore the low-$k_{x}$ limit, while leaving much of the intermediate band governed by the same kernel shape.
At this juncture, we remark that the AEM is a statistical and non-dynamic model, and thus prototypical eddies cannot be interpreted as literal material structures, even if they provide useful building blocks for kinematic prediction. Nevertheless, the inverse-kernel viewpoint provides a bridge between statistical requirements (what kernels are needed to match moments) and geometric hypotheses (what minimL eddy templates can realize those kernels under Biot–Savart induction). A good fit of a one-dimensional premultiplied spectrum does not uniquely identify a specific structure. Many different eddy templates can generate similar $k^{-1}$-like behavior once integrated over a scale-invariant population. Therefore, the fact that the Rankine-rod hairpin AEM matches experiments across many quantities of interest and Reynolds numbers is encouraging, but it is not, by itself, a proof that the flow is literally composed of those hairpins. Stronger discrimination can come from tests that 1D spectra smear out: two-dimensional spectra, cross-spectra, coherence, wall-normal and spanwise components, vorticity-related spectra, and conditional structure signatures.
Inverse problems as a pathway forward.
The inverse formulation suggests several immediate extensions: (i) simultaneous inference of multiple kernels using multiple moments and multiple Reynolds numbers with shared regularization and physically motivated constraints, (ii) inference of 2D spectral kernels (in $(k_{x},k_{z})$) rather than only 1D kernels, and (iii) incorporation of non-Poisson population effects (packets, clustering, exclusion) through pair-correlation corrections to Eq. (11) and their spectral counterparts.
Acknowledgement This research was supported by NASA grant # 80NSSC18M0149 (Technical monitor: Dr. Gary Coleman).
Appendix A: Numerical convergence evidence for the square hairpin + image.
To verify the decay/convergence directly, we performed a Biot–Savart calculation for a simple “square hairpin” U-shape (three straight segments with unit circulation) together with an opposite-sign mirror image across $y=0$, consistent with the construction described around Fig. 5. A small vortex-core regularization was used to smooth the near-filament singularity; this does not affect the far-field decay that controls convergence of the planar integrals.
Far-field decay check. At $\eta=y/h=0.5$, evaluating along $(\xi,0,\eta)$ with $\xi\gg 1$ gives
$u_{1}(10,0,0.5)$ $\approx-3.92\times 10^{-4},$
$u_{1}(20,0,0.5)$ $\approx-4.44\times 10^{-5},$
$u_{1}(40,0,0.5)$ $\approx-5.26\times 10^{-6},$
$u_{1}(80,0,0.5)$ $\approx-6.39\times 10^{-7}.$
Moreover, $u_{1}(\xi,0,0.5)\,\xi^{3}$ is approximately constant (between $\approx-0.33$ and $\approx-0.39$), indicating the scaling
$u_{1}(\rho,\eta)\sim\frac{C(\eta)}{\rho^{3}}\qquad(\rho\to\infty),$ (22)
so that $\big|I_{1}(\eta)-I_{1}(\eta;R)\big|=\mathcal{O}(1/R)$ as predicted by (14) with $p=3$.
Direct growth of $I_{1}$ with domain size. For the same hairpin at $\eta=0.5$, integrating over $R,R]^{2}$ yields: The observed approach is consistent with the $\mathcal{O}(1/R)$ tail implied by (22). In particular, extrapolating with the model $I_{1}(\eta;R)=I_{1,\infty}(\eta)+C/R$ gives
$I_{1,\infty}(\eta=0.5)\approx-0.997.$ (23)
Above the eddy in this toy geometry, the convergence is rapid; e.g. at $\eta=1.2$, $I_{1}(1.2;R)$ approaches 0 quickly (with $I_{1}\approx 0.005$ by $R=24$), again consistent with a well-defined infinite-domain limit.
Practical implication for the homogenization argument.
In the attached-eddy homogenization step, taking $L/h$ “large enough” should be interpreted as choosing $R=L/h$ such that the residual tail $\big|I_{1}(\eta)-I_{1}(\eta;R)\big|$ (and similarly for $I_{ij}$) is negligible at the level of accuracy desired. For the line-vortex eddies considered here, this condition is satisfied because the image construction yields a sufficiently rapid far-field decay of the induced streamwise velocity, ensuring that the planar influence integrals are well-defined in the limit $R\to\infty$.
Table 1: Convergence of the planar influence integral $I_{1}(\eta;R)$ for the square hairpin + image at $\eta=0.5$.
$R=L/h$ $I_{1}(\eta=0.5;R)$
4 0.885-0.885
8 0.941-0.941
12 0.960-0.960
16 0.969-0.969
20 0.975-0.975
24 0.979-0.979